### v2

# Wiedemann-Franz law violation domain for graphene and nonrelativistic systems

Abstract
Systematic and comparative research on Lorenz ratios for graphene and nonrelativistic systems has been studied to identify their Wiedemann-Franz law violation domain. Fermi energy and temperature are the main governing parameters for deciding the values of the Lorenz ratio, which is basically thermal conductivity divided by electrical conductivity times temperature times Lorenz number. Metals as three-dimensional nonrelativistic electron gas locate at higher Fermi energy by temperature domain, where Lorenz ratio remains one. Hence, they obey the Wiedemann-Franz law. By creating higher doping in a two-dimensional graphene system, one can again reach a higher Fermi energy by temperature domain and get a constant Lorenz ratio. For both graphene and nonrelativistic systems, the Lorenz ratio goes below one if we go lower Fermi energy by temperature domain, which is possible for the graphene system by decreasing the doping concentration. Experimentally observed greater than one Lorenz ratio in this lower Fermi energy by temperature domain or Dirac fluid domain indicates that non-fluid expressions of Lorenz ratio should be replaced by fluid-type expressions. We have noticed a divergent trend of Lorenz ratio in the Dirac fluid domain using its fluid-type expression, and it matches with the trend of experimental data.
I Introduction
In 1853, Gustav Wiedemann and Rudolph Franz experimentally discovered a universal or constant value of the ratio between thermal ($\kappa$) and electrical ($\sigma$) conductivity, which are approximately followed by all metals. Later, in 1872, Ludvig Lorenz theoretically realized that the ratio in terms of the universal constants $k_{B}$ (Boltzmann constant) and $e$ (electric charge) is:
$\frac{\kappa}{\sigma T}$ $=$ $\frac{\pi^{2}}{3}\left(\frac{k_{B}}{e}\right)^{2}$ (1)
$=$ $L_{0}=2.445\times 10^{-8}watt\frac{\Omega}{K^{2}}~{},$
where $L_{0}$ is known as the Lorenz number. The fact is very well known from our text book Refs. [1, 2, 3, 4, 5]. In natural unit, $\hbar=c=k_{B}=1$ and $e^{2}=\frac{4\pi}{137}$, we can express temperature $T$ in eV, $\sigma$ in eV and $\kappa$ in eV2. So Lorenz number will be a dimensionless ratio $L_{0}=\frac{137\pi}{12}\approx 35.84$. In the present paper, we will mostly use the natural unit methodology for convenience, but we may go to the actual unit in some cases (whenever necessary).
The Wiedemann-Franz (WF) law has proven remarkably robust for many metallic systems, where electrons are the main electric and thermal charge transportation carriers. Due to the similar mechanism of transportation in free electron theory [1, 2, 3, 4, 5], the dimensionless (in natural unit) ratio of two transport coefficients - thermal and electrical conductivity becomes constant. However, deviations have been observed in many systems like $Li_{0.9}Mo_{6}O_{17}$ [6], $CeCoIn_{5}$ [7], $ZrZn_{2}$ [8], $YbRh_{2}Si_{2}$ [9], $(Pr,Ce)_{2}CuO_{4}$ [10] and $VO_{2}$ [11]. For our convenience, if we define the ratio of thermal and electrical conductivity as $L=\frac{\kappa}{\sigma T}$, and if we call $L/L_{0}$ as Lorenz ratio, then the validity of the WF law will be established from the relations $L=L_{0}$ or $\frac{L}{L_{0}}=1$. On the other hand, relations $\frac{L}{L_{0}}>1$ or $\frac{L}{L_{0}}<1$ reflect violation of the WF law. A large diverging outcomes $\frac{L}{L_{0}}>10^{4}$ is expected in one-dimensional exotic Tomonaga-Luttinger liquids [6]. On the other hand, a strong downward violation of the WF Law $L<L_{0}$ was observed in $WP_{2}$ semimetal [12, 13] and MoP binary compound [14]. This downward violation is also observed in heavy fermion metals [7], marginal ferromagnetic metal [8], anti-ferromagnetic metal [9] and copper-oxide superconducting material [10] at low-temperature regions as well as in metallic vanadium dioxide [11] at high-temperature range ($240$-$340^{\circ}$). Understanding the WF law violation mechanism of that wide range of systems is a non-trivial task and challenge for the theoretical community to explain.
Recently, Crossno et al. [15] have found a similar kind of the WF law violation in graphene systems by tuning the doping concentration and temperature. Unlike standard metal, the Fermi level of graphene can be moved upwards or downwards from the relative Dirac point depending upon the n- or p-type doping of graphene. Fermi energy $(\epsilon_{F})$ or a chemical potential $(\mu)$ at Dirac point is considered zero for undoped graphene and, via n- or p-type doping, will shift it towards the positive or negative direction. The net charge carrier concentration will be tuned from zero to non-zero when one experimentally goes from undoped to doped graphene cases, and theoretically, one goes from $\epsilon_{F}=0$ to $\epsilon_{F}\neq 0$ cases. This kind of tuning possibility of $\epsilon_{F}$ can not be expected for metal systems as it almost remains constant. Their typical values remain within the range $\epsilon_{F}=2-10$ eV, for which one can expect the limit $\epsilon_{F}/T>>1$ at room temperature $T\approx 0.025$ eV. In this limit, one can consider electrons in metal as Fermi gas and use Sommerfeld expansion, which provides a linear T dependent on electron-specific heat. Fermi gas is completely based on the crude non-interacting assumption, but there is a theory of interacting Fermions system, which is popularly known as the Fermi Liquid (FL) Theory. Originally, Landau [16, 17, 18] proposed this phenomenological theory for studying 3He. In FL prescription, interaction is taken care of via the effective mass of electrons by assuming a quasi-particle picture. Hence, some mass-independent quantities like $L$ remain the same as $L_{0}$ for both FL and Fermi gas prescriptions. So, if we define $\epsilon_{F}/T>1$ as the FL domain and $\epsilon_{F}/T>>1$ as the Fermi gas domain, then electron-doped graphene systems almost follow WF law. When we go towards undoped or clean graphene systems with $\epsilon_{F}/T<1$, Fermi liquid theory becomes invalid (as it is commonly accepted). It is concluded from the experimental observation of WF law violation in this $\epsilon_{F}/T<1$ or $\epsilon_{F}/T<<1$ domains, popularly called Dirac fluid (DF) or Dirac Liquid (DL) domain. Theoretical works [19, 20, 21, 22, 23] are attempted to explain this WF law violation in the DF domain. In this regards, electron hydrodynamics (eHD) or fluid dynamics in graphene system, recently observed by Refs [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37], is thought to be linked with this WF law violation. In this context, the present work is planned to explore a transition from non-fluid to fluid type expressions of Lorenz ratio for a 2-dimensional (2D) graphene (G) system with linear energy-momentum relation ($E\propto p$) as well as a 3-dimensional (3D) metal system with nonrelativistic (NR) energy-momentum relation ($E\propto p^{2}$). For mathematical completeness, we have demonstrated all possible cases like 2D-NR, 3D-NR, 2D-G, and 3D-G but our final destination is kept to show how fluid expressions of 2D-G case can be associated with WF law violation and how non-fluid expressions of 3D-NR case, followed by metal systems, obey WF law. In the direction of eHD, our earlier comparative research on ideal [38] and viscous dissipating [39] parts of energy-momentum tensor calculation clearly demonstrate that the graphene hydrodynamics is neither relativistic hydrodynamics nor nonrelativistic hydrodynamics. The present work has gone through similar comparative research on electrical conductivity, thermal conductivity, and Lorenz ratio for a systematic search of the WF law violation domain.
The article is organized as follows, in Sec. (II), the formalism part of WF law calculations for different cases. The results and discussions are in Sec. (III) and summarised by Sec.( IV).
II Formalism
Graphene is a 2-dimensional single atomic layer of carbon atoms tightly bound into a honeycomb lattice [40]. Near the Dirac point, electrons in graphene follow the linear dispersion relation
$\epsilon\left(k\right)=pv_{F},$ (2)
where $v_{F}$ is the Fermi velocity of electrons in graphene, whose values remain within the range $0.003c$-$0.01c$ [41]. In one direction, electrons in graphene do not follow a quadratic relation between energy and momentum ($\epsilon=p^{2}/(2m)$) like in the nonrelativistic case. On the other hand, its velocity ($v_{F}$) is not close to the speed of light ($c$), so no relativistic corrections are expected. In that sense, electrons in graphene follow neither relativistic nor nonrelativistic. So, its different properties may not be the same as expected for traditional nonrelativistic or relativistic matters as shown in Refs. [38, 39], for thermodynamics [38] and dissipative viscous [39] properties. In this regard, thermal and electrical conductivity for the graphene (G) case may also be interesting to compare with nonrelativistic (NR) and ultrarelativistic (UR) cases, which are studied in the present work. In sub-sec. (II.1), we have gone through the first non-fluid type framework description, which is traditionally adopted in standard books of solid state physics [1, 2, 3, 4, 5]. Next, in sub-sec. (II.2), we have extended our framework towards fluid type description.
II.1 Non-fluid description
For systematic study and understanding, we will calculate the Lorenz ratios for the 3D system and then the 2D system. The NR, UR, and G case expressions of the Lorenz ratio will be addressed step by step for each system. Then, we will also examine the limiting conditions using Sommerfeld expansion.
3D-NR: Let us first discuss the 3D-NR case. The well-known Drude’s formula for the electrical conductivity of nonrelativistic electrons in 3D solids (metals) is
$\sigma_{NR}^{3D}=\frac{ne^{2}\tau_{c}}{m}=\frac{ne^{2}\lambda}{mv_{F}},$ (3)
and the thermal conductivity is
$\kappa_{NR}^{3D}=\frac{1}{3}nv_{F}\lambda\,\prescript{3D}{NR}{[C_{V}]_{e}},$ (4)
where, $n$ is the number (electron) density, $\tau_{c}$ is the relaxation time, $m$ is electron mass, $\lambda=v_{F}\tau_{c}$ is the mean free path, and $[C_{V}]_{e}$ is the electronic specific heat per particle. After taking the ratio of thermal and electrical conductivity, we get
$\frac{\kappa_{NR}^{3D}}{\sigma_{NR}^{3D}}=\frac{2}{3}\frac{\epsilon_{F}}{e^{2}}\,\prescript{3D}{NR}{[C_{V}]_{e}},$ (5)
where $\epsilon_{F}=\mu=\frac{1}{2}mv_{F}^{2}$ is the Fermi energy.
Here, in Eq. (5), $[C_{V}]_{e}$ is the main important thing to calculate the conductivity ratio. We will use the two definitions of $[C_{V}]_{e}$:
  1. 1.
    $[C_{V}]_{e1}=\frac{\partial}{\partial T}\Big{(}\frac{U}{N}\Big{)}\Bigg{|}_{V,\ \epsilon_{F}}$,
  2. 2.
    $[C_{V}]_{e2}=\frac{1}{N}\frac{\partial U}{\partial T}\Bigg{|}_{V,\ \epsilon_{F}}$.
Here, for subsequent simplicity, we will use the notations like NR1 and G1 concerning $[C_{V}]_{e1}$ and NR2 and G2 dealing with $[C_{V}]_{e2}$. The former definition prescribes taking $T$ derivative of internal energy per particle $u=U/N$, while the latter definition says to take $T$ derivative of internal energy $U$ and then normalized by $N$. The detailed calculation can be seen in Appendix C. Let us first put the former specific heat,
[$\prescript{3D}{NR1}{[C_{V}]_{e1}}=\frac{3}{2}k_{B}\Bigg{[}\frac{5}{2}\frac{f_{\frac{5}{2}}\left(A\right)}{f_{\frac{3}{2}}\left(A\right)}-\frac{3}{2}\frac{f_{\frac{3}{2}}\left(A\right)}{f_{\frac{1}{2}}\left(A\right)}\Bigg{]},$ (6)
in Eq. (5), the Lorenz ratio for 3D-NR will be
$\frac{L_{NR1}^{3D}}{L_{0}}=\frac{3}{\pi^{2}}\frac{\epsilon_{F}}{k_{B}T}\Bigg{[}\frac{5}{2}\frac{f_{\frac{5}{2}}\left(A\right)}{f_{\frac{3}{2}}\left(A\right)}-\frac{3}{2}\frac{f_{\frac{3}{2}}\left(A\right)}{f_{\frac{1}{2}}\left(A\right)}\Bigg{]}.$ (7)
Here,
[…]
is the standard Fermi integral function (see details in Appendix B) with $A=e^{\epsilon_{F}/k_{B}T}$. Using Sommerfeld lemma, in the limit of $\epsilon_{F}/k_{B}T=lnA>>1$, the electronic specific heat becomes
[$\prescript{3D}{NR1}{[C_{V}]_{e1}}=\frac{\pi^{2}}{2}\frac{k_{B}^{2}T}{\epsilon_{F}},$ (9)
and then Eq. (7) becomes
$L_{NR1}^{3D}=\frac{\pi^{2}}{3}\left(\frac{k_{B}}{e}\right)^{2}=L_{0},$ (10)
which is the so-called Wiedemann-Franz law for metals.
Next, using another definition of $[C_{V}]_{e2}$ for the 3D-NR case, whose general expression will be (see Appendix C)
[$\prescript{3D}{NR2}{[C_{V}]_{e2}}=\frac{3}{2}k_{B}\Bigg{[}\frac{5}{2}\frac{f_{\frac{5}{2}}\left(A\right)}{f_{\frac{3}{2}}\left(A\right)}-\frac{\epsilon_{F}}{k_{B}T}\Bigg{]}~{},$ (11)
we will get the Lorenz ratio as
$\frac{L_{NR2}^{3D}}{L_{0}}=\frac{3}{\pi^{2}}\frac{\epsilon_{F}}{k_{B}T}\Bigg{[}\frac{5}{2}\frac{f_{\frac{5}{2}}\left(A\right)}{f_{\frac{3}{2}}\left(A\right)}-\frac{\epsilon_{F}}{k_{B}T}\Bigg{]}.$ (12)
In the limit of $\epsilon_{F}/T=lnA>>1$, Sommerfeld expansion of Eq. (11) becomes
[$\prescript{3D}{NR2}{[C_{V}]_{e2}}=\frac{3\pi^{2}}{4}\frac{k_{B}^{2}T}{\epsilon_{F}},$ (13)
and then the Eq. (12) becomes
$L_{NR2}^{3D}=\frac{\pi^{2}}{2}\left(\frac{k_{B}}{e}\right)^{2}=1.5L_{0}.$ (14)
Reader can see that Eq. (10) and (14), both are independent of $T$ , $\epsilon_{F}$ and are made by universal constants like $k_{B}$ and $e$. So, the domain of $T-\epsilon_{F}$ plane, where $\frac{L}{L_{0}}$ or $L$ becomes independent of $T$ , $\epsilon_{F}$ , can be considered as Wiedemann-Franz law obeying domain and remaining will be Wiedemann-Franz law violating domain. Although, in terms of $L_{0}$ , $L$ is $1.5$ factor times larger for Eq. (14).
3D-G: Now, we consider a hypothetical 3D system, which follows graphene-type dispersion relation $\epsilon=v_{F}p$. In this case, the electrical conductivity
$\sigma_{G}^{3D}=\frac{ne^{2}\tau_{c}}{\epsilon_{F}}v_{F}^{2}=\frac{ne^{2}\lambda}{\epsilon_{F}}v_{F},$ (15)
and the thermal conductivity,
$\kappa_{G}^{3D}=\frac{1}{3}nv_{F}\lambda\prescript{3D}{G}{[C_{V}]_{e}},$ (16)
will form the ratio
$\frac{\kappa_{G}^{3D}}{\sigma_{G}^{3D}}=\frac{1}{3}\frac{\epsilon_{F}}{e^{2}}\prescript{3D}{G}{[C_{V}]_{e}},$ (17)
where the relation between the Fermi energy and the Fermi momentum will be $\epsilon_{F}=p_{F}v_{F}$ for graphene.
Using two possible expressions (see Appendix C) of specific heat for 3D-G system,
[$\prescript{3D}{G1}{[C_{V}]_{e1}}=3k_{B}\Bigg{[}4\frac{f_{4}\left(A\right)}{f_{3}\left(A\right)}-3\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}\Bigg{]}$ (18)
and
[$\prescript{3D}{G2}{[C_{V}]_{e2}}=3k_{B}\Bigg{[}4\frac{f_{4}\left(A\right)}{f_{3}\left(A\right)}-\frac{\epsilon_{F}}{k_{B}T}\Bigg{]},$ (19)
in Eq. (17), we get the Lorenz ratios
$\frac{L_{G1}^{3D}}{L_{0}}=\frac{\kappa_{G1}^{3D}}{\sigma_{G1}^{3D}TL_{0}}=\frac{3}{\pi^{2}}\frac{\epsilon_{F}}{k_{B}T}\Bigg{[}4\frac{f_{4}\left(A\right)}{f_{3}\left(A\right)}-3\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}\Bigg{]}$ (20)
and
$\frac{L_{G2}^{3D}}{L_{0}}=\frac{\kappa_{G2}^{3D}}{\sigma_{G2}^{3D}TL_{0}}=\frac{3}{\pi^{2}}\frac{\epsilon_{F}}{k_{B}T}\Bigg{[}4\frac{f_{4}\left(A\right)}{f_{3}\left(A\right)}-\frac{\epsilon_{F}}{k_{B}T}\Bigg{]}$ (21)
respectively. In the Sommerfeld limit, Eqs. (18) and (19) will be converted to
[$\prescript{3D}{G1}{[C_{V}]_{e1}}=\pi^{2}\frac{k_{B}^{2}T}{\epsilon_{F}},$ (22)
and
[$\prescript{3D}{G2}{[C_{V}]_{e2}}=3\pi^{2}\frac{k_{B}^{2}T}{\epsilon_{F}},$ (23)
and so, Eqs. (20) and (21) become
$L_{G1}^{3D}=\frac{\pi^{2}}{3}\left(\frac{k_{B}}{e}\right)^{2}=L_{0},$ (24)
and
$L_{G2}^{3D}=\pi^{2}\left(\frac{k_{B}}{e}\right)^{2}=3L_{0},$ (25)
respectively.
2D-NR: Now, let us go for 2D cases. In the case of 2D nonrelativistic system, the ratio of thermal and electrical conductivity can be written as
$\frac{\kappa_{NR}^{2D}}{\sigma_{NR}^{2D}}=\frac{\epsilon_{F}}{e^{2}}\prescript{2D}{NR}{[C_{V}]_{e}}.$ (26)
Using two possible expressions (see Appendix C) of specific heat for 2D-NR system,
[$\prescript{2D}{NR1}{[C_{V}]_{e1}}=k_{B}\Bigg{[}2\frac{f_{2}\left(A\right)}{f_{1}\left(A\right)}-\frac{f_{1}\left(A\right)}{f_{0}\left(A\right)}\Bigg{]},$ (27)
and
[$\prescript{2D}{NR2}{[C_{V}]_{e2}}=k_{B}\Bigg{[}2\frac{f_{2}\left(A\right)}{f_{1}\left(A\right)}-\frac{\epsilon_{F}}{k_{B}T}\Bigg{]},$ (28)
in Eq. (26), we get the Lorenz ratios
$\frac{L_{NR1}^{2D}}{L_{0}}=\frac{\kappa_{NR1}^{2D}}{\sigma_{NR1}^{2D}TL_{0}}=\frac{3}{\pi^{2}}\frac{\epsilon_{F}}{k_{B}T}\Bigg{[}2\frac{f_{2}\left(A\right)}{f_{1}\left(A\right)}-\frac{f_{1}\left(A\right)}{f_{0}\left(A\right)}\Bigg{]}$ (29)
and
$\frac{L_{NR2}^{2D}}{L_{0}}=\frac{\kappa_{NR2}^{2D}}{\sigma_{NR2}^{2D}TL_{0}}=\frac{3}{\pi^{2}}\frac{\epsilon_{F}}{k_{B}T}\Bigg{[}2\frac{f_{2}\left(A\right)}{f_{1}\left(A\right)}-\frac{\epsilon_{F}}{k_{B}T}\Bigg{]}$ (30)
respectively. In the Sommerfeld limit (SL), Eqs. (29) and (30) will be converted to
$L_{NR1}^{2D}=\frac{\pi^{2}}{3}\left(\frac{k_{B}}{e}\right)^{2}=L_{0},$ (31)
$L_{NR2}^{2D}=\frac{\pi^{2}}{3}\left(\frac{k_{B}}{e}\right)^{2}=L_{0}.$ (32)
2D-G: Now, we are taking an actual 2D graphene case. The ratio of $\kappa_{G}^{2D}$ and $\sigma_{G}^{2D}$ is given by
$\frac{\kappa_{G}^{2D}}{\sigma_{G}^{2D}}=\frac{1}{2}\frac{\epsilon_{F}}{e^{2}}\prescript{2D}{G}{[C_{V}]_{e}}.$ (33)
Using two possible expressions (see Appendix C) of specific heat for 2D-G system,
[$\prescript{2D}{G1}{[C_{V}]_{e1}}=2k_{B}\Bigg{[}3\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}-2\frac{f_{2}\left(A\right)}{f_{1}\left(A\right)}\Bigg{]},$ (34)
and
[$\prescript{2D}{G2}{[C_{V}]_{e2}}=2k_{B}\Bigg{[}3\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}-\frac{\epsilon_{F}}{k_{B}T}\Bigg{]},$ (35)
in Eq. (33), we get the Lorenz ratios
$\frac{L_{G1}^{2D}}{L_{0}}=\frac{\kappa_{G1}^{2D}}{\sigma_{G1}^{2D}TL_{0}}=\frac{3}{\pi^{2}}\frac{\epsilon_{F}}{k_{B}T}\Bigg{[}3\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}-2\frac{f_{2}\left(A\right)}{f_{1}\left(A\right)}\Bigg{]}$ (36)
and
$\frac{L_{G2}^{2D}}{L_{0}}=\frac{\kappa_{G2}^{2D}}{\sigma_{G2}^{2D}TL_{0}}=\frac{3}{\pi^{2}}\frac{\epsilon_{F}}{k_{B}T}\Bigg{[}3\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}-\frac{\epsilon_{F}}{k_{B}T}\Bigg{]}$ (37)
respectively. The SL of Eqs. (36) and (37) will be
$L_{G1}^{2D}=\frac{\pi^{2}}{3}\left(\frac{k_{B}}{e}\right)^{2}=L_{0},$ (38)
and
$L_{G2}^{2D}=\frac{2\pi^{2}}{3}\left(\frac{k_{B}}{e}\right)^{2}=2L_{0}$ (39)
respectively.
3D-UR and 2D-UR: The expressions of electrical and thermal conductivity for 3D-UR and 2D-UR will be the same as those expressions for 3D-G and 2D-G if $v_{F}$ is replaced by the speed of light $c$. Readers can notice that specific heat and Lorenz ratios for 3D-G and 2D-G are independent of $v_{F}$, so those expressions can also be used for 3D-UR and 2D-UR systems. Since our focal quantity is the Lorenz ratio in the results, we will not discuss it further. The expressions of 3D-UR and 2D-UR systems coincide with those of 3D-G and 2D-G systems.
II.2 An approach towards the fluid description
The previous subsection provides the expression of $\kappa$ and $\sigma$ using a standard solid state physics framework, where no fluid concept has been entered. So, we can call that description a non-fluid description. In the present section, we will try to build an approach towards a fluid description by using relaxation time approximation-based Boltzmann transport equations. Here, we are using the terminology ”towards the fluid description” because the actual phenomenological fluid description [21] for explaining the Wiedemann-Franz law violation in the experimental 2D graphene system is quite rigorous. Aim of the present work is to zoom in the intermediate journey from non-fluid to fluid description. Interestingly, we will get a Wiedemann-Franz law violating domain in $T-\epsilon_{F}$ plane for all cases. Let us first address the brief calculations of the Lorenz ratio for 2D-G systems; then, we will write down its final expressions for other possible systems like 2D-NR, 3D-NR, and 3D-G instead of repeating similar calculations.
2D-G: Let us assume a local thermalization picture of electron fluid in a 2D-G system, where the equilibrium distribution function (see Appendix A)
$f_{0}\left(T(x^{\mu})\right)=\frac{1}{e^{\left(\epsilon-\epsilon_{F}\right)/k_{B}T(x^{\mu})}+1},$ (40)
will be $x^{\mu}$-dependent due to temperature profile $T(x^{\mu})$. Here, $x^{\mu}=(ct,x^{i})$, a four-dimensional coordinate is considered for general notation, but we have to take care of $i=1,2$ for the 2D system and $i=1,2,3$ for the 3D system.
Now, let us first write down the macroscopic definitions for electrical and thermal conductivity
${\vec{J}}$ $=$ $\sigma{\vec{E}},$
${\vec{Q}}$ $=$ $\kappa{\vec{\nabla}}T~{},$ (41)
where electrical current vector ${\vec{J}}$ and heat flow vector ${\vec{Q}}$ can have microscopic expressions:
${J_{i}}$ $=$ $ge\int\frac{d^{2}p}{h^{2}}v_{i}\delta f_{\sigma},$
${Q_{i}}$ $=$ $g\int\frac{d^{2}p}{h^{2}}\epsilon v_{i}\delta f_{\kappa}.$ (42)
Here, we are assuming that the external electric field $E_{i}$ and the temperature gradient ${\vec{\nabla}}T$ will create deviations $\delta f_{\sigma}$ and $\delta f_{\kappa}$ respectively from the equilibrium distribution $f_{0}$.
It is relaxation time approximation (RTA) based on Boltzmann’s transport equation (BTE) [42]
$\frac{\partial f}{\partial t}+\frac{\partial x^{i}}{\partial t}\frac{\partial f}{\partial x^{i}}+\frac{\partial p^{i}}{\partial t}\frac{\partial f}{\partial p^{i}}=-\frac{\delta f}{\tau_{c}}~{},$ (43)
which will guide us to guess appropriate form of $\delta f_{\sigma}$ and $\delta f_{\kappa}$. Considering $f=f_{0}+\delta f\approx f_{0}$ in left hand side of Eq. (43) and local thermalization assumption of $f_{0}$, given in Eq. (40), we can simplify Eq. (43) as
$v^{i}\frac{\partial f_{0}}{\partial x^{i}}+eE^{i}\frac{\partial f_{0}}{\partial p^{i}}$ $=$ $-\frac{{\delta f}_{\sigma}}{\tau_{c}}-\frac{{\delta f}_{\kappa}}{\tau_{c}}$
$v^{i}\frac{\partial T}{\partial x^{i}}\frac{\partial\epsilon}{\partial T}\frac{\partial f_{0}}{\partial\epsilon}+eE^{i}\frac{\partial\epsilon}{\partial p^{i}}\frac{\partial f_{0}}{\partial\epsilon}$ $=$ $-\frac{{\delta f}_{\sigma}}{\tau_{c}}-\frac{{\delta f}_{\kappa}}{\tau_{c}}$
$v^{i}\frac{\partial T}{\partial x^{i}}[C_{V}]_{e}\frac{\partial f_{0}}{\partial\epsilon}+eE^{i}v_{i}\frac{\partial f_{0}}{\partial\epsilon}$ $=$ $-\frac{{\delta f}_{\sigma}}{\tau_{c}}-\frac{{\delta f}_{\kappa}}{\tau_{c}}.$ (44)
Here, we consider the approximation $[C_{V}]_{e}\approx\frac{\partial\epsilon}{\partial T}$, and one can expect again two possible definitions of specific heat as discussed in the earlier section. From Eq. (44), we can get the form of $\delta f_{\sigma}$ and $\delta f_{\kappa}$ as
$\delta f_{\sigma}$ $=$ $eE^{i}v_{i}\left(-\frac{\partial f_{0}}{\partial\epsilon}\right)\tau_{c},$
$\delta f_{\kappa}$ $=$ $v^{i}\left(-\frac{\partial f_{0}}{\partial\epsilon}\right)[C_{V}]_{e}\left(\frac{\partial T}{\partial x^{i}}\right)\tau_{c},$ (45)
with
$\left(-\frac{\partial f_{0}}{\partial\epsilon}\right)=\beta f_{0}(1-f_{0}).$ (46)
Using Eq. (45) in Eq. (42) and then comparing with Eq. (41), we will get the final expressions of electrical and thermal conductivity:
$\sigma$ $=$ $ge^{2}\frac{v_{F}^{2}}{2}\tau_{c}\int\frac{d^{2}p}{h^{2}}\beta f_{0}(1-f_{0})$
$\implies\sigma_{G}^{2D}$ $=$ $2\pi k_{B}\tau_{c}\left(\frac{e}{h}\right)^{2}f_{1}\left(A\right)T,$ (47)
and
$\kappa$ $=$ $g\epsilon\frac{v_{F}^{2}}{2}\tau_{c}[C_{V}]_{e}\int\frac{d^{2}p}{h^{2}}\beta f_{0}(1-f_{0})$
$\implies\kappa_{G}^{2D}$ $=$ $\frac{4\pi k_{B}^{2}\tau_{c}}{h^{2}}[C_{V}]_{e}\,f_{2}\left(A\right)T^{2}.$ (48)
Now, using two different forms of specific heat, given in Eqs. (34) and (35) in Eq. (48), we get
$\kappa_{G1}^{2D}=\frac{8\pi k_{B}^{3}\tau_{c}}{h^{2}}\Bigg{[}3\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}-2\frac{f_{2}\left(A\right)}{f_{1}\left(A\right)}\Bigg{]}f_{2}\left(A\right)T^{2},$ (49)
and
$\kappa_{G2}^{2D}=\frac{8\pi k_{B}^{3}\tau_{c}}{h^{2}}\Bigg{[}3\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}-\frac{\epsilon_{F}}{k_{B}T}\Bigg{]}f_{2}\left(A\right)T^{2},$ (50)
and hence, corresponding Lorenz ratios will be
$\frac{L_{G1}^{2D}}{L_{0}}=\frac{12}{\pi^{2}}\Bigg{[}3\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}-2\frac{f_{2}\left(A\right)}{f_{1}\left(A\right)}\Bigg{]}\frac{f_{2}\left(A\right)}{f_{1}\left(A\right)},$ (51)
and
$\frac{L_{G2}^{2D}}{L_{0}}=\frac{12}{\pi^{2}}\Bigg{[}4\frac{f_{4}\left(A\right)}{f_{3}\left(A\right)}-\frac{\epsilon_{F}}{k_{B}T}\Bigg{]}\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}$ (52)
respectively.
2D-NR After doing the same calculation as the previous one for a 2-dimensional nonrelativistic system, the final general expressions of the electrical and two possible thermal conductivities and their corresponding Lorenz ratios $L_{NR}^{2D}/L_{0}$ are given by as below:
$\sigma_{NR}^{2D}=4\pi k_{B}\tau_{c}\left(\frac{e}{h}\right)^{2}f_{1}\left(A\right)T,$ (53)
$\kappa_{NR1}^{2D}=\frac{8\pi k_{B}^{3}\tau_{c}}{h^{2}}\Bigg{[}2\frac{f_{2}\left(A\right)}{f_{1}\left(A\right)}-\frac{f_{1}\left(A\right)}{f_{0}\left(A\right)}\Bigg{]}f_{2}\left(A\right)T^{2},$ (54)
$\kappa_{NR2}^{2D}=\frac{8\pi k_{B}^{3}\tau_{c}}{h^{2}}\Bigg{[}2\frac{f_{2}\left(A\right)}{f_{1}\left(A\right)}-\frac{\epsilon_{F}}{k_{B}T}\Bigg{]}f_{2}\left(A\right)T^{2}~{},$ (55)
$\frac{L_{NR1}^{2D}}{L_{0}}=\frac{6}{\pi^{2}}\Bigg{[}2\frac{f_{2}\left(A\right)}{f_{1}\left(A\right)}-\frac{f_{1}\left(A\right)}{f_{0}\left(A\right)}\Bigg{]}\frac{f_{2}\left(A\right)}{f_{1}\left(A\right)},$ (56)
$\frac{L_{NR2}^{2D}}{L_{0}}=\frac{6}{\pi^{2}}\Bigg{[}2\frac{f_{2}\left(A\right)}{f_{1}\left(A\right)}-\frac{\epsilon_{F}}{k_{B}T}\Bigg{]}\frac{f_{2}\left(A\right)}{f_{1}\left(A\right)}.$ (57)
3D-G: For the 3D-G case, the expression of the electrical conductivity and two possible thermal conductivities and their corresponding Lorenz ratios $L_{G}^{3D}/L_{0}$ are given as below:
$\sigma_{G}^{3D}=16\pi k_{B}^{2}\tau_{c}\left(\frac{e^{2}}{3h^{3}v_{F}}\right)f_{2}\left(A\right)T^{2},$ (58)
$\kappa_{G1}^{3D}=\frac{48\pi k_{B}^{4}\tau_{c}}{h^{3}v_{F}}\Bigg{[}4\frac{f_{4}\left(A\right)}{f_{3}\left(A\right)}-3\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}\Bigg{]}f_{3}\left(A\right)T^{3},$ (59)
$\kappa_{G2}^{3D}=\frac{48\pi k_{B}^{4}\tau_{c}}{h^{3}v_{F}}\left[4\frac{f_{4}\left(A\right)}{f_{3}\left(A\right)}-\frac{\epsilon_{F}}{k_{B}T}\right]f_{3}\left(A\right)T^{3},$ (60)
$\frac{L_{G1}^{3D}}{L_{0}}=\frac{27}{\pi^{2}}\Bigg{[}4\frac{f_{4}\left(A\right)}{f_{3}\left(A\right)}-3\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}\Bigg{]}\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)},$ (61)
$\frac{L_{G2}^{3D}}{L_{0}}=\frac{27}{\pi^{2}}\Bigg{[}4\frac{f_{4}\left(A\right)}{f_{3}\left(A\right)}-\frac{\epsilon_{F}}{k_{B}T}\Bigg{]}\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}.$ (62)
3D-NR: Now again, the same type of calculation is done for a 3-dimensional nonrelativistic system; we get all expressions of the electrical, thermal conductivity, and the ratio $L_{NR}^{3D}/L_{0}$ in a more general form, which can be expressed as
$\sigma_{NR}^{3D}=\frac{2e^{2}k_{B}^{\frac{3}{2}}\tau_{c}}{m}\left(\frac{2\pi m}{h^{2}}\right)^{\frac{3}{2}}f_{\frac{3}{2}}\left(A\right)T^{\frac{3}{2}}~{},$ (63)
$\kappa_{NR1}^{3D}=\frac{15k_{B}^{\frac{7}{2}}\tau_{c}}{2m}\left(\frac{2\pi m}{h^{2}}\right)^{\frac{3}{2}}\Bigg{[}\frac{5}{2}\frac{f_{\frac{5}{2}}\left(A\right)}{f_{\frac{3}{2}}\left(A\right)}-\frac{3}{2}\frac{f_{\frac{3}{2}}\left(A\right)}{f_{\frac{1}{2}}\left(A\right)}\Bigg{]}$
$f_{\frac{5}{2}}\left(A\right)T^{\frac{5}{2}},$ (64)
$\kappa_{NR2}^{3D}=\frac{15k_{B}^{\frac{7}{2}}\tau_{c}}{2m}\left(\frac{2\pi m}{h^{2}}\right)^{\frac{3}{2}}\Bigg{[}\frac{5}{2}\frac{f_{\frac{5}{2}}\left(A\right)}{f_{\frac{3}{2}}\left(A\right)}-\frac{\epsilon_{F}}{k_{B}T}\Bigg{]}$
$f_{\frac{5}{2}}\left(A\right)T^{\frac{5}{2}},$ (65)
$\frac{L_{NR1}^{3D}}{L_{0}}=\frac{45}{4\pi^{2}}\Bigg{[}\frac{5}{2}\frac{f_{\frac{5}{2}}\left(A\right)}{f_{\frac{3}{2}}\left(A\right)}-\frac{3}{2}\frac{f_{\frac{3}{2}}\left(A\right)}{f_{\frac{1}{2}}\left(A\right)}\Bigg{]}\frac{f_{\frac{5}{2}}\left(A\right)}{f_{\frac{3}{2}}\left(A\right)}~{},$ (66)
$\frac{L_{NR2}^{3D}}{L_{0}}=\frac{45}{4\pi^{2}}\Bigg{[}\frac{5}{2}\frac{f_{\frac{5}{2}}\left(A\right)}{f_{\frac{3}{2}}\left(A\right)}-\frac{\epsilon_{F}}{k_{B}T}\Bigg{]}\frac{f_{\frac{5}{2}}\left(A\right)}{f_{\frac{3}{2}}\left(A\right)}.$ (67)
III Results
Figure 1: The Lorenz ratio regarding the chemical potential for non-fluid descriptions of graphene case (a) for 3D and (b) for 2D expressions.
Figure 2: The Lorenz ratio regarding the chemical potential for non-fluid descriptions of nonrelativistic case (a) for 3D and (b) for 2D expressions.
In the above section, we have calculated all the general expressions of electrical and thermal conductivity and their ratios to check the validity of the Wiedemann-Franz law for all possible systems like 2D and 3D graphene (G) and nonrelativistic (NR) cases. Using those final expressions, the present section is intended to go with their numerical estimations. During the result generation, we will use natural unit $k_{B}=\hbar=c=1$ for our convenience. Dimensionless quantities Lorenz ratio ($L/L_{0}$) and $\epsilon_{F}/T$ will be taken as the y and x axes of all graphs.
Figure 3: Left: Lorenz ratio of 2D-G systems by using fluid-type expression. Right: Comparison of specific heat and fluid aspect in 2D-G system.
Figure 4: Left: The Lorenz ratio of fluid and non-fluid for 2D-G system. Right: The Lorenz ratio of the mixture of fluid and non-fluid with fitting parameters for 2D-G system.
In Fig. 1 and Fig. 2, we explored the Lorenz ratio ($\frac{L}{L_{0}}$) by using non-fluid (NF) type expressions. Fig. 1(a) and (b) have shown the Lorenz ratio versus $\frac{\epsilon_{F}}{T}$ for 3D-G and 2D-G systems respectively. By using the non-fluid (NF) type of Eq. (20) and its Sommerfeld limit Eq. (24), we plotted the black solid and dotted lines in Fig. 1(a). We used the specific heat ${[C_{V}]_{e1}}=\frac{\partial}{\partial T}\Big{(}\frac{U}{N}\Big{)}$, which is traditionally used in solid state or non-fluid systems. We are getting $L/L_{0}=1$ in the limit of $\epsilon_{F}/T\gg 1$ or Fermi Liquid (FL) regime of graphene case, which is as expected for most of the metal cases. The deviation of the Lorenz ratio from 1 has been started below $\frac{\epsilon_{F}}{T}\approx 10$ as we notice from the black solid line. We have also plotted the red solid line by using Eq. (21), where another definition of specific heat ${[C_{V}]_{e2}}$ is used in the thermal conductivity expression. In the case of SL of 3D-G (the red dotted line), the Lorenz ratio is three instead of one. For the general case (the red solid line), the Lorenz ratio deviates from this constant line $3$ below $\frac{\epsilon_{F}}{T}\approx 10$. For the 2D-G system, similar kinds of black and red solid lines are plotted in Fig. 1(b) by using Eq. (36) and Eq. (37). Their respective SLs by using Eq. (38) and Eq. (39), are plotted as black and red horizontal dotted lines. We considered the surface area $(S)$ in place of volume $(V)$ in the 2D system. The Lorenz ratio when we take ${[C_{V}]_{e2}}$, the ratio becomes 2 (the red-dotted line).
The Fig. 2(a) and (b), similar to Fig. 1, display the Lorenz ratio in terms of $\frac{\epsilon_{F}}{T}$ for non-fluid expressions of NR case. Though the real metals have a constant chemical potential within the range, $\epsilon_{F}=2-10$ eV, we imagine that it can be tunable from 00 to $\infty$. In Fig. 2(a), we notice that the WF law is valid in the SL of the 3D metal case, expressed by Eq. (10) (green dot-dashed line). We have taken $T=60^{\circ}$ K ($\approx 0.005$ eV), so metal range $\epsilon_{F}=2$-$10$ eV will give the ratio $\frac{\epsilon_{F}}{T}=400$-$2000$ ($\gg 1$), which is quite good domain for SL. Therefore, being situated in the SL domain, metal always follows the WF law. A slight larger value $L=1.5L_{0}$ comes from Eq. (14) (blue dotted line), when specific heat ${[C_{V}]_{e2}}$ is taken. A similar kind of behavior can be seen in Fig. 2(b) for 2D-NR systems, but interestingly, SLs in both cases, given by Eqs. (31) and (32) are coincided at $L=L_{0}$ (blue-dotted and green dot-dashed lines). It probably indicates that if one builds 2D metal systems, then it will also follow the WF law as their Fermi energy range is again expected to be situated in the SL domain.
Before going to the next graphical results, let us briefly analyze what we have learned from Fig. 1(a), (b) and Fig. 2(a), (b). Grossly, the non-fluid type expressions of Lorenz ratios for all cases - 3D-NR, 2D-NR, 3D-G, and 2D-G are saturated near 1 in the SL range, which means that the WF law is obeyed in the SL range. Among all four cases, 3D-NR and 2D-G are realistic cases. The former is applicable to metal systems, whose Fermi energies remain constant and within the SL range, so they always follow the WF law. Latter case - the non-fluid expressions of Lorenz ratio for 2D-G can be applicable for realistic graphene system with high doping concentration ($\epsilon_{F}$ is quite large or $\epsilon_{F}/T\gg 1$), where the WF law is obeyed well again [15]. However, by decreasing the doping concentration, when the graphene system approaches the charge-neutral points ($\epsilon_{F}$ is quite small or $\epsilon_{F}/T\ll 1$), the WF law violation domain is observed [15]. According to our non-fluid expressions, the WF law violation is also expected in the $\epsilon_{F}/T\ll 1$ domain but the outcome $L/L_{0}<1$ is not matching with experimental outcome $L/L_{0}>1$ by Crossno et al. [15]. A possible explanation of $L/L_{0}>1$ in the appearance of electron hydrodynamics [21] property in the domain $\epsilon_{F}/T<1$, which is popularly called the Dirac fluid domain. Present work is firstly intended to highlight the overlooked possibility of Wiedemann-Franz law violation of any systems like NR, G or UR even we consider the non-fluid expression of $\kappa$ and $\sigma$. Then, its next attempt is to show the transition from non-fluid (NF) to fluid (F) pictures in a systematic way. An approach towards the fluid description is addressed in formalism part Sec: (II.2). Let us come to the numerical estimation for fluid-type formalism.
For numerical curves of fluid-type expressions, let us focus only on the 2D-G case. Although formalism of other cases with an Appendix is provided in the present article, one may go for result generation for graphene case. Using Eqs. (51) and (52), black and red solid lines are plotted in the left panel of Fig. 3. They are saturated at the values of 2 and 4, respectively, in the high doping or $\epsilon_{F}/T\gg 1$ domain. It means that if the fluid type expression is valid in this high doping domain 2D-G system, then the WF law may not be obeyed. However, the experimental data [15] claims that the WF law is well obeyed in the high doping or $\epsilon_{F}/T\gg 1$ domain. This means that the non-fluid expression of the Lorenz ratio should be applicable to this domain instead of fluid-type expression. The SL of Eqs. (51) and (52) will also be
$\frac{L_{G1}^{2D}}{L_{0}}=2+\frac{2\pi^{2}}{3}\frac{k_{B}^{2}T^{2}}{\epsilon_{F}^{2}},$ (68)
$\frac{L_{G2}^{2D}}{L_{0}}=4+\frac{4\pi^{2}}{3}\frac{k_{B}^{2}T^{2}}{\epsilon_{F}^{2}},$ (69)
which are plotted by black and red dotted lines, respectively, in the left panel of Fig. 3. They blow up from their saturated values (2 and 4) when $\epsilon_{F}/T$ or doping concentration decreases. The experimental data of Crossno et al. [15] is pasted, where S1 and S2 represent the samples of the 2D-G system in terms of lower to higher values of minimum charge carrier density. Theoretically, zero charge carrier density can be reached at $\epsilon_{F}/T\rightarrow 0$ but experimentally, a minimum charge carrier density [15] will be achieved. So, from Fig.3(a), we notice that our fluid-based expression of $\frac{L_{G2}^{2D}}{L_{0}}$, given in Eq.(51), (52), can’t able to explain the experimental data points. It indicates that our formalism needs next level modification. Due to this step by step development of our formalism from (standard) non-fluid to (new directional) fluid description, we put the name of Sec: (II.2) as ”towards the fluid description”. Now if we revisit our RTA based kinetic theory formalism, then we can find that we have systematically considered the local thermalization concept of fluid. In this concept, we consider fluid as a large number fluid elements, having velocity $u^{\mu}$, temperature T and chemical potential $\epsilon_{F}$. Now, fluid elements will be at different position and time $(x^{\mu})$, so, considering $u^{\mu}(x^{\mu})$, $T(x^{\mu})$, $\epsilon_{F}(x^{\mu})$ as function of $x^{\mu}$ can map mathematically the local thermalization concept. Since present work is only focus on thermal and electrical conductivity, so only $T(x^{\mu})$ is considered and others $u^{\mu}(x^{\mu})$ and $\epsilon_{F}(x^{\mu})$ are ignored. Reader can see this consideration in Eq.(44), where the information of temperature gradient $\frac{\partial T}{\partial x^{i}}$ is entered. Now the another term specific heat $[C_{V}]_{e}$ $\approx\frac{\partial u}{\partial T}$, which is basically considered as temperature derivative of average energy of electron. This average energy of electron information can be modified into enthalpy per electron $h=\frac{u+P}{n}$ in fluid picture. Another point is that electron-hole plasma picture will be important when we go $\frac{\epsilon_{F}}{T}<1$ domain (towards charge neutrality point). So, we should take enthalpy per net electron $h=\frac{u+P}{n}$, where $u=u_{e}+u_{h}$, $P=P_{e}+P_{h}$ and $n=n_{e}-n_{h}$ will be respectively total internal energy density, pressure and net charge density. We have adopted the replacement
$[C_{V}]_{e}\approx\frac{\partial u}{\partial T}\,\rightarrow\,\frac{h}{T}=\frac{u+P}{nT}$ (70)
and using it in Eq. (48), we will get a better fluid-type expression of thermal conductivity:
$\kappa_{G}^{2D}=\frac{4\pi k_{B}^{2}\tau_{c}}{h^{2}}\left(\frac{u+P}{nT}\right)\,f_{2}\left(A\right)T^{2}$ (71)
and Lorenz ratio:
$\left(\frac{L_{G1}^{2D}}{L_{0}}\right)=\frac{6}{\pi^{2}}\left(\frac{u+P}{nT}\right)\frac{f_{2}\left(A\right)}{f_{1}\left(A\right)}~{}.$ (72)
For comparison between two quantities $C_{V}^{e}$ and $\frac{h}{T}=\frac{u+P}{nT}$, we have plotted them in the right panel of Fig.3(b). Next, in Fig.4(a), we have drawn dashed-line by using Eq.(72). Here, one can notice that the Lorenz ratio shows a divergence trend in $\frac{\epsilon_{F}}{T}<1$ domain, which in favor of experimental data $S_{1}$ and $S_{2}$. We have also plotted non-fluid(NF) expression of Lorenz ratio, given in Eq.(36), which is also in favor of experimental data in $\frac{\epsilon_{F}}{T}>1$ domain. Based on this data favoring theoretical curves, we have classified three basic regions: $(1)$ Fluid (F) domain, $(2)$ Non-Fluid (NF) domain and $(3)$ Mixed or (F+NF) domain, which are discussed one by one (below).
  • Fluid (F) domain: Fluid based Lorenz ratio, given in Eq.(72) valid well with experimental data in the range $0.70.2<\frac{\epsilon_{F}}{T}<$ 0.7.
  • Non-Fluid (NF) domain: Non-Fluid based Lorenz ratio, given in Eq.(36), valid well with experimental data in the ranges (a) $102.5<\frac{\epsilon_{F}}{T}<$ 10 and (b) $10<\frac{\epsilon_{F}}{T}<\infty$. We can called subrange (b) as ohmic non-fluid or NF(Ohmic) domain, since we get WF law or $\frac{L}{L_{0}}=1$ for this $\frac{\epsilon_{F}}{T}>10$ domain, where Fermi-Dirac distribution function can be assumed as step function. However, for subrange (a), we get $\frac{L}{L_{0}}<1$ because assumption of step function is not further valid here. Corresponding domains are also marked in Fig.3(b), where we can notice that $C_{V}^{e}\propto\frac{T}{\epsilon_{F}}$ follows in NF(Ohmic) domain (due to validation of electron step function like distribution) but a transition from $C_{V}^{e}\propto\frac{T}{\epsilon_{F}}$ to $C_{V}^{e}=2k_{B}\Bigg{[}3\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}-2\frac{f_{2}\left(A\right)}{f_{1}\left(A\right)}\Bigg{]}$ is observed in NF domain ($102.5<\frac{\epsilon_{F}}{T}<$ 10). Reader can understand this transition from Eq.(34).
  • Mixed or (F+NF) domain: The domain $2.50.7<\frac{\epsilon_{F}}{T}<$ 2.5 is marked as mixed or (F+NF) domain as we have to go with replacement $C_{V}^{e}\rightarrow\frac{u+P}{nT}$ or transit from NF based quantity $C_{V}^{e}$ to fluid-based quantity $\frac{u+P}{nT}$. This domain may be considered as a grey zone and we may put an open question for further research- How this transition happens? Interestingly, reader can notice a $\frac{\epsilon_{F}}{T}$ independent trend approximately of $C_{V}^{e}$ and $\frac{u+P}{nT}$ within this (F+NF) domain.
If we analyze the F domain of Fig.3(b) and Fig.4(a), then we can identify that the divergent trend of $\frac{u+P}{nT}$ and $\frac{L}{L_{0}}$ are inter-related. Net density $n$ trends to zero when $\frac{\epsilon_{F}}{T}$ trends to zero and that is why $\frac{L}{L_{0}}\propto\frac{1}{n}$ trends to infinite. So, we may connect the divergent trend of Lorenz ratio in Dirac fluid domain ($\frac{\epsilon_{F}}{T}<1$) with the fluid properties of electron-hole plasma in graphene, where fluid quantity enthalpy per particle will be the main resource for that divergence.
To make a bridge between the F and NF domain, we have used two switching functions.
$a_{F}\left(x\right)=\frac{1}{e^{b_{F}\left(x-x_{F}\right)}+1}~{},$ (73)
$a_{NF}\left(x\right)=\frac{1}{e^{b_{NF}\left(x-x_{NF}\right)}+1}~{}.$ (74)
where $x=\frac{\epsilon_{F}}{T}$ and $b_{F}$, $b_{NF}$, $x_{F}$, $x_{NF}$ are tuning parameters. Using these switching functions, we have prescribed a sandwich expression of Lorenz ratio as
$\frac{L}{L_{0}}=\left(\frac{L}{L_{0}}\right)_{F}\times a_{F}+\left(\frac{L}{L_{0}}\right)_{NF}\times a_{NF}~{},$ (75)
where NF and F Lorenz ratio can be written from Eq. (51) and (72) in a compact form.
$\left(\frac{L}{L_{0}}\right)_{F/NF}=\frac{6}{\pi^{2}}\,\big{[}C_{V}^{e}\big{]}_{F/NF}\,\frac{f_{2}\left(A\right)}{f_{1}\left(A\right)}\,(A=e^{x})~{},$ (76)
with
$\big{[}C_{V}^{e}\big{]}_{F}=\frac{u+P}{nT}~{},$ (77)
$\big{[}C_{V}^{e}\big{]}_{NF}=2k_{B}\Bigg{[}3\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}-2\frac{f_{2}\left(A\right)}{f_{1}\left(A\right)}\Bigg{]}~{}.$ (78)
Using Eq.(76), we have plotted blue and red solid curves to fit experimental data point $S_{1}$(blue circles) and $S_{1}$(red circles). For blue line, tuning parameters are taken as $x_{F}=0.2$, $x_{NF}=0.8$, $b_{F}=2$, $b_{NF}=0.7$ and for red line, tuning parameters are taken as $x_{F}=0.9$, $x_{NF}=2$, $b_{F}=2$, $b_{NF}=0.7$. By using a better version of x-dependent switching functions, one may go with better fitting of experimental data. However, instead of quantitative exact fitting, the core aim of the present work may be considered as its qualitative message or understanding on the F and NF domain and their transition.
This qualitative message of transition from NF to F is in agreement with earlier theoretical works by Lucas et al. [21]. Though final expressions of Lorenz ratio of Ref. [21] and the present work are little different but both works have similar qualitative message- NF to F transition is associated with the WF law violation and enhancement of Lorenz ratio more than one. There are other works also which provide little different reasons for getting WF law violations. Ref. [20] has linked this fact with presence of band gap and bipolar diffusion, while Ref. [19] has built thermodynamical expression of electron-hole plasma in terms of polylogarithm function, which can provide a greater than one values of Lorenz ratio. So, many more theoretical research on this topic are expected in future for getting some converging understanding.
IV Summary and Interpretation
In summary, we have gone through comparative research on Lorenz ratios for 3D, 2D NR and G systems for a systematic search of the WF law violation domain. During our analysis, we found the WF law violation domain of Fermi energy (normalized by a fixed temperature), where the Lorenz ratio deviates from one. First, we have used standard solid state book-based non-fluid expressions of Lorenz ratio for the 3D-NR case, which are also applied for 2D-NR, 3D-G, and 2D-G cases. For all cases, the Lorenz ratio remains constant in the higher values of normalized Fermi energy. This fact is well in agreement with the universal validity of the WF law in metal systems, which can be described by the 3D-NR case of the electron gas having larger Fermi energy ($2-10$ eV) or normalized Fermi energy 400-2000 at $60^{\circ}$ K. Depending on the definition of specific heat at constant volume and Fermi energy, we will get two possible saturated values of Lorenz ratios - 1 and 2, where former is experimentally observed (average) value but one may mark the saturating domain as the WF law obeying domain. If one identifies two domains - less and greater than $10$ value of normalized Fermi energy (i.e. $\frac{\epsilon_{F}}{T}<10$ and $\frac{\epsilon_{F}}{T}>10$), then the former domain of all cases show the deviation of WF law. This tuning of normalized Fermi energy is conveniently possible for 2D-G system only via tuning the electron doping. Recently measured Lorenz ratio by Crossno et al. [15] indicates the violation of the WF law in the lower normalized Fermi energy domain, which is popularly called as Dirac fluid (DF) domain. Using standard solid state or non-fluid theory for 2D-G system, we also find WF law violation in the lower normalized Fermi energy domain. However, non-fluid theory gets less than one values of Lorenz ratio, while experiment gets its greater than one values. It demands that we may have to transit towards a fluid-type description from the standard non-fluid expressions of the Lorenz ratio. So far, to the best of our knowledge, this straightforward thermodynamical phase space analysis of Lorenz ratio, based on non-fluid theory is missing in the literature. Present work is fulfilling this gap, which may be considered one of the important contents of present article. A next level content of present article is its fluid based calculations, whose outcomes are summarized in the next paragraph.
Towards the fluid description, we have approached in two steps. In the first step, we have used Boltzmann transport equation for 2D-G system, where local thermalization concept is mainly implimented for fluid description. If we compare the NF and F expressions of Lorenz ratios, then we can find that only via specific heat, Fermi integral functions are entered into the NF expression of Lorenz ratio. On the other hand, F expressions of electric and thermal conductivity and hence Lorenz ratio carry additional Fermi integral functions, which are completely appeared because of local thermalization aspect of fluid description. During this first step attempt of fluid description, we get a little deviated outcomes of Lorenz ratio with respect to NF results. Sommerfeld limit of F-based Lorenz ratios give a divergent outcomes instead of constant values (it was found in its NF-based expression). This divergent trend of Lorenz ratio in lower normalized Fermi energy is experimentally observed, so this first step fluid description results may be considered as a preliminary hints of NF to F transition in the Dirac fluid domain or lower Fermi energy domain. Observing the quantitative mismatch between experimental data and first step fluid theory, we have proceeded to second step extension of fluid theory. Here, we have gone through the replacement of specific heat by enthalpy per net electron (divided by temperature). This replacement is based on the intuition for transition from quasi-particle concept to fluid concept. In NF picture, electron average energy is the quasi-particle concept and its temperature derivative gives specific heat concept. Whereas, enthalpy per net electron is the fluid concept for electron-hole plasma in Dirac fluid domain. Interestingly, we notice that this conceptual transformation and replacement provide an enhancement of Lorenz ratio in Dirac fluid domain and their theoretical values are quite close to experimental data points.
After developing this step by step NF to F frameworks, we have identified four regions along the axis of normalized Fermi energy -
  • Above $10$: Lorenz ratio is one and it can be called hardcore NF Ohmic domain. Sommerfeld limit of our NF expression can explain it easily.
  • Within $2.5$ to $10$: Lorenz raio become less than one and it can be called NF but non-ohmic type domain. Our NF expression of Lorenz ratio and experimental data both hint this domain.
  • Within $0.7$ to $2.5$: It is the transition domain from NF to F by showing less than one to greater than one values of Lorenz ratio.
  • Below $0.7$: Here, Lorenz ratio becomes greater than one and it can be called Fluid domain as enhancing trend in fluid theory and experimental data towards lower doping range is noticed.
At the end, for final quantitative data matching, we have demonstrated a sandwich estimation of F and NF results. However, present article intend to draw the attention of reader on the qualitative message of NF to F transition. This association of enhancement Lorenz ratio observation in the Dirac fluid domain with the NF to F transition is agreement with earlier theoretical works. Although, some alternative explanations are also addressed by some earlier works, more theoretical studies are probably required for better understanding about this WF law violation in graphene system.
Acknowledgements.
This work was partly (T.Z.W. and C.W.A.) supported by the Doctoral Fellowship in India (DIA) program of the Ministry of Education, Government of India. The authors thank the other members of eHD club Sesha P. Vempati, Ashutosh Dwibedi, Narayan Prasad, Bharat Kukkar, and Subhalaxmi Nayak.
Appendix A FERMI-DIRAC DISTRIBUTION FUNCTION
The Fermi-Dirac distribution function is
$f_{0}\left(\epsilon\right)=\frac{1}{e^{\beta\left(\epsilon-\epsilon_{F}\right)}+1}=\frac{1}{A^{-1}e^{\beta\epsilon}+1},$ (79)
where $\epsilon$ = energy of fermions and $\epsilon_{F}$ = fermi energy, and $A$ is the fugacity of the system described by
$A=\exp{\left({\frac{\epsilon_{F}}{k_{B}T}}\right)}.$
And the derivative of the distribution function with respect to energy is
$-\frac{\partial f_{0}}{\partial\epsilon}=\frac{\beta e^{\beta\left(\epsilon-\epsilon_{F}\right)}}{\left(e^{\beta\left(\epsilon-\epsilon_{F}\right)}+1\right)^{2}}=\frac{\partial}{\partial\epsilon_{F}}\left(\frac{1}{e^{\beta\left(\epsilon-\epsilon_{F}\right)}+1}\right).$ (80)
Appendix B FERMI-DIRAC FUNCTION
  • The Fermi-Dirac Function
[…]
    where $f_{\nu}(A)$ is known as the Fermi-Dirac function.
    Case.1. When $A$ is small, then the Fermi-Dirac function can be written in a series form which is
    $f_{\nu}(A)=$ $A-\frac{A^{2}}{2^{\nu}}+\frac{A^{3}}{3^{\nu}}-\frac{A^{4}}{4^{\nu}}+\dots$ (82)
    $=\sum_{n=1}^{\infty}\left(-1\right)^{n-1}\frac{A^{n}}{n^{\nu}}.$ (83)
    Case.2. When $A$ is too much small $\left(\epsilon_{F}<<k_{B}T\right)$, then the function becomes simplified as $A$
    $f_{\nu}(A)=A,$ (84)
    but if we take $\epsilon_{F}=0$, the function becomes unity.
    Case.3. When the temperature is very small, and the Fermi energy has some finite value, then the Fermi-Dirac function can be written according to Sommerfeld lemma, which gives the expression of the function as
    $f_{\nu}\left(A\right)=\frac{\alpha^{\nu}}{\Gamma\left(\nu+1\right)}\Bigg{[}1+\nu\left(\nu-1\right)\frac{\pi^{2}}{6}\frac{1}{\alpha^{2}}+\nu\left(\nu-1\right)\left(\nu-2\right)\left(\nu-3\right)\frac{7\pi^{4}}{360}\frac{1}{\alpha^{4}}+\dots\Bigg{]},$ (85)
    where $\alpha$ is given by
    $\alpha=\ln{A}=\frac{\epsilon_{F}}{k_{B}T}~{}.$
    And
    $\Gamma\left(\nu+1\right)=\nu!~{},$
    so, using this lemma, we can calculate the value of the Fermi-Dirac function for different values of $\nu$ for this particular type of case.
    Case.4. When $\left(\epsilon_{F}>>k_{B}T\right)$, then the function can be written only using the zeroth order term of Sommerfeld lemma expression.
    The derivative of Fermi function with respect to Fermi energy
    $\frac{\partial f_{\nu}\left(A\right)}{\partial\epsilon_{F}}=\beta f_{\nu-1}\left(A\right).$ (86)
    The derivative of Fermi function with respect to temperature
    $\frac{\partial f_{\nu}\left(A\right)}{\partial T}=\frac{1}{A}f_{\nu-1}\left(A\right)\frac{\partial A}{\partial T}.$ (87)
    The one identity for the function can be written as
    $\frac{1}{A}\frac{\partial A}{\partial T}=-\epsilon_{F}\beta^{2}k_{B}=-\frac{\epsilon_{F}}{k_{B}T^{2}}.$ (88)
Appendix C ELECTRONIC SPECIFIC HEAT
For 2D system graphene, which follows the linear dispersion relation, with the help of the density of state method, we can calculate the number density, total internal energy, and the electronic specific heat.
So, the number of energy states in energy range $\epsilon$ to $\epsilon+d\epsilon$ and the surface area $S$ is written as
$D\left(\epsilon\right)d\epsilon=g\frac{S}{\left(2\pi\hbar\right)^{2}}\frac{2\pi}{v_{F}^{2}}\epsilon d\epsilon.$ (89)
Now, the total number of particles at any value of temperature can be calculated as
[…]
After plugging the value of $D\left(\epsilon\right)d\epsilon$ in the above equation, we get
[…]
After solving this integration, we get the final, more general expression of number density, which is given by
$N=g\frac{2\pi S}{h^{2}v_{F}^{2}}\frac{\Gamma\left(2\right)}{\beta^{2}}f_{2}\left(A\right)$ (92)
$\implies n=\frac{k_{B}^{2}}{\pi\hbar^{2}v_{F}^{2}}f_{2}\left(A\right)T^{2}~{}.$ (93)
The total internal energy of a system can be calculated as
[…]
By substituting the value of $D\left(\epsilon\right)d\epsilon$ in the above equation and solving the integration, we get the final, more general expression of total internal energy, which is given by
$U=g\frac{2\pi S}{h^{2}v_{F}^{2}}\frac{\Gamma\left(3\right)}{\beta^{3}}f_{3}\left(A\right)$ (95)
$\implies u=\frac{U}{S}=g\frac{2\pi}{h^{2}v_{F}^{2}}\frac{\Gamma\left(3\right)}{\beta^{3}}f_{3}\left(A\right).$ (96)
Now, from the above equations, we get
$\frac{u}{n}=\frac{U}{N}=2k_{B}T\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}.$ (97)
Let us define the specific heat capacity of the electron by taking the temperature derivative of internal energy per electron ($U/N=u/n$) while keeping surface area $S$ and $\epsilon_{F}$ as constants:
[$\prescript{2D}{G1}{[C_{V}]_{e1}}=\Bigg{[}\frac{\partial}{\partial T}\Big{(}\frac{U}{N}\Big{)}\Bigg{]}_{S,\ \epsilon_{F}}.$ (98)
Using the Eq. (97), the specific heat is
[$\prescript{2D}{G1}{[C_{V}]_{e1}}=2k_{B}\Bigg{[}\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}+T\frac{\partial}{\partial T}\left(\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}\right)\Bigg{]}.$ (99)
Now, the derivative part after using the identity (87) of the Fermi-Dirac function can be written as
$\frac{\partial}{\partial T}\left(\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}\right)$ $=\frac{1}{f_{2}\left(A\right)}\frac{\partial}{\partial T}\left(f_{3}\left(A\right)\right)-f_{3}\left(A\right)\frac{\partial}{\partial T}\left(\frac{1}{f_{2}\left(A\right)}\right)$
$=\frac{1}{A}\frac{f_{2}\left(A\right)}{f_{2}\left(A\right)}\frac{\partial A}{\partial T}-\frac{f_{3}\left(A\right)}{\Big{[}f_{2}\left(A\right)\Big{]}^{2}}\frac{\partial}{\partial T}f_{2}\left(A\right)$
$=\frac{1}{A}\frac{\partial A}{\partial T}-\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}\frac{f_{1}\left(A\right)}{f_{2}\left(A\right)}\frac{1}{A}\frac{\partial A}{\partial T}.$
The final form, we get
$\frac{\partial}{\partial T}\left(\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}\right)=\frac{2}{T}\Bigg{[}\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}-\frac{f_{2}\left(A\right)}{f_{1}\left(A\right)}\Bigg{]}.$ (100)
After substituting the value from equation (100) into equation (99), we get a more general form of electronic specific heat given by
[$\prescript{2D}{G1}{[C_{V}]_{e1}}=2k_{B}\Bigg{[}3\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}-2\frac{f_{2}\left(A\right)}{f_{1}\left(A\right)}\Bigg{]}~{}.$ (101)
Let us go for another definition of specific heat per electron
[$\prescript{2D}{G2}{[C_{V}]_{e2}}=\frac{1}{N}\frac{\partial U}{\partial T}\Bigg{|}_{S,\ \epsilon_{F}}.$ (102)
From Eq. (95), the total internal energy is
$\prescript{2D}{G}{U}=\gamma T^{3}f_{3}\left(A\right),$
where
$\gamma=g\frac{2\pi S}{h^{2}v_{F}^{2}}\Gamma\left(3\right)k_{B}^{3}.$
Then, we will take the derivative of just the above equation to $T$ at constant Fermi energy $\epsilon_{F}$ and surface area $S$; we get
$\frac{\partial U}{\partial T}\Big{|}_{S,\epsilon_{F}}$ $=\gamma\Bigg{[}3T^{2}f_{3}\left(A\right)+T^{3}f_{2}\left(A\right)\frac{1}{A}\frac{\partial A}{\partial T}\Bigg{]}$ (103)
$=\gamma T^{2}f_{2}\left(A\right)\Bigg{[}3\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}+T\frac{1}{A}\frac{\partial A}{\partial T}\Bigg{]}~{}.$ (104)
So,
[$\prescript{2D}{G2}{[C_{V}]_{e2}}=2k_{B}\Bigg{[}3\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}-\frac{\epsilon_{F}}{k_{B}T}\Bigg{]}.$ (105)
By doing a similar type of calculation for 2D-NR case, we will get two different specific heat expressions as
[$\prescript{2D}{NR1}{[C_{V}]_{e1}}=k_{B}\Bigg{[}2\frac{f_{2}\left(A\right)}{f_{1}\left(A\right)}-\frac{f_{1}\left(A\right)}{f_{0}\left(A\right)}\Bigg{]},$ (106)
and
[$\prescript{2D}{NR2}{[C_{V}]_{e2}}=k_{B}\Bigg{[}2\frac{f_{2}\left(A\right)}{f_{1}\left(A\right)}-\frac{\epsilon_{F}}{k_{B}T}\Bigg{]}.$ (107)
Next, for 3D-NR and 3D-G cases also, one can repeat the calculations and find the expressions of specific heat:
[$\prescript{3D}{NR1}{[C_{V}]_{e1}}=\frac{3}{2}k_{B}\Bigg{[}\frac{5}{2}\frac{f_{\frac{5}{2}}\left(A\right)}{f_{\frac{3}{2}}\left(A\right)}-\frac{3}{2}\frac{f_{\frac{3}{2}}\left(A\right)}{f_{\frac{1}{2}}\left(A\right)}\Bigg{]},$ (108)
[$\prescript{3D}{NR2}{[C_{V}]_{e2}}=\frac{3}{2}k_{B}\Bigg{[}\frac{5}{2}\frac{f_{\frac{5}{2}}\left(A\right)}{f_{\frac{3}{2}}\left(A\right)}-\frac{\epsilon_{F}}{k_{B}T}\Bigg{]},$ (109)
[$\prescript{3D}{G1}{[C_{V}]_{e1}}=3k_{B}\Bigg{[}4\frac{f_{4}\left(A\right)}{f_{3}\left(A\right)}-3\frac{f_{3}\left(A\right)}{f_{2}\left(A\right)}\Bigg{]},$ (110)
[$\prescript{3D}{G2}{[C_{V}]_{e2}}=3k_{B}\Bigg{[}4\frac{f_{4}\left(A\right)}{f_{3}\left(A\right)}-\frac{\epsilon_{F}}{k_{B}T}\Bigg{]}.$ (111)